A companys flexible budget for the range of 30,000 units to 36,000 units of production showed variable overhead costs of $3.60 per unit and fixed overhead costs of $70,000. The company incurred total overhead costs of $180,100 while operating at a volume of 40,000 units. The total controllable cost variance is Multiple Choice $2,100 unfavorable. $2100 favorable. 33.900 favorable. $6,000 favorable $33.900 unfavorable
Hassock Corp. produces woven wall hangings. It takes 4 hours of direct labor to produce a single wall hanging. Hassocks standard labor cost is $16 per hour. During August, Hassock produced 17.900 units and used 72.190 hours of direct labor at a total cost of $1152,040. What is Hassocks labor efficiency variance for August? Multiple Choice 12.440 unfavorable $6.000 favorable 9,440 unfavorable 9440 favorable 6440 untavorable
A company provided the following direct materials cost information. Compute the total direct materials cost variance Standard costs assigned: Direct materials standard cost (405,000 units e $2.00/unit) 810,000 Actual costs: Direct Materials costs incurred (403,750 units $2.20/unit) $888,250 Multiple Choice
Actual costs: Direct Materials costs incurred (403,750 units $2.20/unit) $888,250 Multiple Choice $2,500 Favorable. $80,750 Unfavorable. $80,750 Favorable $78.250 Favorable. $78.250 Unfavorable.
Answer #1

1) Total Controllable Cost Variance = Actual Cost- Standard Cost

= 180,100- (40,000x3.6)+70,000

=33,900 (Favorable)

2)Labor Efficiency Variance = SR(AH-SH)

=16(72190-(17900X4))

=9,440 (Unfavorable)

3)Total Direct Materials Cost Variance = Actual Cost- Standard Cost

=888,250 - 810,000

=78,150 (Unfavorable)